OpenStudy (anonymous):

(will reward with thank you and medal) Graph G(x)=x^3. Write the equation for F(x) if the graph is flipped over the x-axis.

OpenStudy (anonymous):

Go through the process if you want to!

OpenStudy (anonymous):

The equation y = f(x) + c shifts the graph of y = f(x) up c units. ... negative on the outside of a function flips the graph with respect to the x−axis.) ... shift left; if we subtract a constant inside a function, we shift right. ... (d) f(x) = x3; g(x) = 1. 4. (x − 5)3 + 2. Now we have three transformations. ... by a factor of 4 and a flip with respect.

OpenStudy (anonymous):

johnny?

OpenStudy (anonymous):

Hi, was just reading and was a bit confused. Is that all there is to it?

OpenStudy (anonymous):

yeah but if you do not get the equation I could get you a link.

OpenStudy (anonymous):

A link would be nice ..

OpenStudy (anonymous):

okay let me get one

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Also what's with 1.4.?

OpenStudy (anonymous):

If I may ask

OpenStudy (anonymous):

Your welcome, oh and it is one of the percents, by when you do the equation

OpenStudy (anonymous):

So what you're saying is that g(x) = 1.4(x − 5)3 + 2? Or is there more steps after that? Please bare with me I'm not good at Math.

OpenStudy (anonymous):

Yes, g(x) = 1.4(x − 5)3 + 2 also I'am a bit shaky in math, I'm a 94 or 90 percent at math.

OpenStudy (anonymous):

Thanks for your help! :)

OpenStudy (anonymous):

Your welcome again

OpenStudy (anonymous):

It gets alot easier in college.

OpenStudy (anonymous):

I hope so haha!